# Single vibronic level fluorescence spectra from Hagedorn wavepacket dynamics

### v1

@@ Abstract
In single vibronic level (SVL) fluorescence experiments, the electronically excited initial state is also excited in one or several vibrational modes. Whereas the time-independent approach of computing all contributing Franck-Condon factors becomes impractical in large systems, a time-dependent formalism has not been applied to simulate emission from arbitrary initial vibrational levels. Here, we apply Hagedorn functions, which are products of a Gaussian and carefully generated polynomials, to represent SVL initial states. Under an at most quadratic potential, the Hagedorn functions are exact solutions to the time-dependent Schrödinger equation and can be propagated using the same equations of motion as a simple Gaussian wavepacket. Having developed an efficient recursive algorithm to compute the overlaps between two Hagedorn wavepackets, we can now evaluate emission spectra from arbitrary vibronic levels using a single trajectory. Here, we use two-dimensional global harmonic models to validate the method by comparing it with quantum split-operator calculations and to demonstrate the effects of displacement, distortion (squeezing), and Duschinsky rotation on SVL spectra. Additionally, we show the practicality of the Hagedorn approach in a high-dimensional system on a displaced, distorted, and Duschinsky-rotated harmonic model with 100 degrees of freedom.
@@ I Introduction
Tuned narrow-band lasers can be used to induce specific vibronic transitions and to prepare a population of molecules excited to a single vibronic level (SVL). The subsequent fluorescence decay of such a system can then be measured and provide important information about excited-state relaxation, higher vibrational structures of the ground state, as well as the vibronic coupling between ground and excited states.[1, 2, 3, 4, 5]
@@
The simulation of SVL spectra of small molecules has been carried out using time-independent sum-over-state expressions,[6, 7, 8, 9, 10, 11, 12] but the computation of Franck-Condon factors in larger molecules becomes challenging with the increasing number of vibrational states involved, especially due to non-separable overlap integrals under Duschinsky mixing effects. A time-dependent approach is more efficient for larger systems and constitutes a more fundamental method that can readily incorporate Duschinsky rotation, Herzberg-Teller,[13, 14, 15] and temperature effects.[16, 17] While time-dependent approaches have been applied to SVL[18, 19] and similar experiments,[20, 21, 22] their practical application has so far been confined to initial states that are singly excited and only in one vibrational mode.
@@
For absorption or emission from the ground vibrational state, the thawed Gaussian approximation[23, 24, 25] (TGA) has been successfully applied to simulate vibrationally resolved electronic spectra in situations where a single Gaussian wavefunction is sufficient to adequately describe the dynamics.[26, 27, 28, 16, 29, 17, 30] However, the SVL emission process involving a vibrationally excited initial state cannot be described by a single Gaussian wavepacket even at the initial time. Within the harmonic approximation, it may instead be described by a Gaussian multiplied by a polynomial at all times. Tapavicza’s generating function approach[19] can treat initial states singly excited in one mode, whereas the extended TGA[31, 14, 32, 33] (ETGA) may accommodate simultaneous single excitations in multiple modes, but only in systems without Duschinsky rotation.iiiWith simultaneous multi-mode excitation, although each polynomial factor is linear in its respective dimension, their tensor product is no longer linear. However, without Duschinsky rotation effects, a system remains fully separable at all times and can be treated as several one-dimensional problems. A more general and robust method is needed to address situations where the simultaneous excitation of multiple vibrational modes to higher levels is involved.
@@
By generalizing Dirac’s ladder operators for one-dimensional quantum harmonic oscillators, Hagedorn introduced a special pair of raising and lowering ladder operators in arbitrary dimensions that generate a series of functions in the form of a Gaussian wavepacket multiplied by a polynomial.[34, 35, 36, 37, 38, 39] Unlike simple direct products of one-dimensional Hermite functions, Hagedorn functions are exact solutions to the time-dependent Schrödinger equation (TDSE) with a general many-dimensional harmonic potential when the parameters of the associated Gaussian are propagated with classical-like equations of motion. The Hagedorn functions also form a complete orthonormal basis that may be used to expand numerically exact arbitrary solutions to the TDSE with a general potential.[40] However, in this work, we consider exclusively Hagedorn wavepackets consisting of only one such Hagedorn function at all times. Henceforth, we will use “Hagedorn wavepackets” and “Hagedorn functions” as synonyms.
@@
Despite their promising properties[38, 41, 42, 43, 39] and a few applications in quantum and chemical physics,[40, 44, 45, 46, 47, 48] Hagedorn wavepackets have yet to be applied to spectroscopic simulation, since one important ingredient, an efficient way to compute the correlation functions between two Hagedorn wavepackets, was missing. Having developed an efficient algebraic algorithm for this overlap in Ref. 49, we can now apply the Hagedorn approach to the simulation of SVL spectra.
@@
Here, we first investigate SVL emissions from different vibrational levels in three two-dimensional, two-state harmonic models, examining the effects on spectra due to nuclear displacement, mode distortion (squeezing), and Duschinsky rotation. To validate our approach, we compare the Hagedorn-wavepacket results to “exact” quantum split-operator calculations and, where applicable, the results from (extended) TGA simulations. Importantly, the propagation of Hagedorn wavepackets does not incur any additional computational cost in harmonic potentials beyond that of propagating a single Gaussian wavepacket. The Hagedorn approach is thus suitable for much higher dimensions than grid-based quantum methods. As a demonstration, we compute spectra in a two-state harmonic model system with 100 vibrational degrees of freedom.
@@ II Theory
Let $|K\rangle\equiv|{e,K}\rangle$ denote a specific vibrational state on the potential energy surface of the excited electronic state $e$, where $K=(K_{1},\dots,K_{D})$ is the multi-index of non-negative integers specifying the vibrational quantum numbers in the $D$ normal modes. Within the Condon approximation, the emission rate from $|K\rangle$ to the ground electronic state $g$ may be obtained by the Fourier transform[50, 13, 19]
@@
Solving the time-dependent Schrödinger equation is a challenging numerical problem. Hagedorn’s semiclassical wavepackets offer a practical approach to evolve a wavepacket in higher-dimensional systems.[34, 37, 39] In Hagedorn’s parametrization,[40, 51] a normalized $D$-dimensional Gaussian wavepacket is written as
$\varphi_{0}[q_{t},p_{t},Q_{t},P_{t},S_{t}](q)=\frac{1}{(\pi\hbar)^{D/4}\sqrt{\det(Q_{t})}}\\ \times\exp\left(\frac{i}{\hbar}\left(\frac{1}{2}x^{T}\cdot P_{t}\cdot Q_{t}^{-1}\cdot x+p_{t}^{T}\cdot x+S_{t}\right)\right],$ (3)
with the shifted position $x:=q-q_{t}$, where $S_{t}$ is a phase factor related to the classical action, $q_{t}$ and $p_{t}$ represent the position and momentum of the wavepacket’s center, and $Q_{t}$ and $P_{t}$ are complex-valued $D$-dimensional matrices related to the position and momentum covariances. The two matrices replace the “width matrix” $A_{t}=P_{t}\cdot Q_{t}^{-1}$ used in Heller’s parametrization[50, 39, 51] and satisfy the symplecticity conditions [38, 39]
@@
Hagedorn defined a special pair of raising and lowering operators,
$A^{\dagger}$ $\equiv A^{\dagger}[q_{t},p_{t},Q_{t},P_{t}]:=\frac{i}{\sqrt{2\hbar}}\left(P_{t}^{\dagger}\cdot\hat{x}-Q_{t}^{\dagger}\cdot\hat{y}\right),$ (6)
$A$ $\equiv A[q_{t},p_{t},Q_{t},P_{t}]\hphantom{{}^{\dagger}}:=-\frac{i}{\sqrt{2\hbar}}\left(P_{t}^{T}\cdot\hat{x}-Q_{t}^{T}\cdot\hat{y}\right),$ (7)
@@
where $\langle j\rangle=(0,\ldots,1,\ldots 0)$ is the $j$-th unit vector in $D$ dimensions.[34, 37, 38, 39] In general, the polynomial factor cannot be expressed as a tensor product of univariate polynomials.[36, 53, 54, 39]
@@
which are the same as the equations for propagating a Gaussian wavepacket.[37, 39, 51] Here, the subscript $t$ denotes the time, $m$ the real symmetric mass matrix (scalar in mass-weighted normal-mode coordinates), and $L_{t}$ is the Lagrangian. Remarkably, with Hagedorn wavepackets, the SVL emission from any vibrational level can be evaluated from the same trajectory $(q_{t},p_{t},Q_{t},P_{t},S_{t})$ at zero cost beyond propagating a simple Gaussian.
To finally obtain SVL spectra from the dynamics results, the evaluation of overlap integrals (2) between two Hagedorn wavepackets with different Gaussian parameters is needed. However, a straightforward analytical expression, like that available for overlaps between two Gaussian wavepackets, has remained elusive for Hagedorn wavepackets. Past applications of Hagedorn wavepackets require numerical quadrature for integrals, which becomes challenging in higher dimensions.[40, 44, 45, 46, 47, 48] In Ref. 49, we have proposed instead a recursive algebraic scheme to compute these overlaps. This allows us to apply Hagedorn wavepackets in much higher dimensions, paving the way for their spectroscopic applications. The cost of computing the overlaps becomes negligible compared to the costly polyatomic electronic structure calculations, which only need to be carried out once for the central Gaussian trajectory, and do not have to be repeated for the vibrationally excited initial states.
Hagedorn wavepackets can be seen as a generalization of coherent states,[55, 37, 41, 54] and our approach is thus connected to the generating function formalism proposed by Huh and Berger.[18] Although their expressions, relying on multivariate Hermite polynomials, are also general for arbitrary initial vibrational levels, no implementation or application was given. As far as we know, the only practical formulation is that of Tapavicza,[19] which is so far limited to single excitation in one mode. The Hagedorn approach, in addition to its generality, also provides a more intuitive time-dependent picture and offers a straightforward propagation scheme.
In this study, we focus solely on harmonic models so that the results can be analyzed without additional approximation or errors. Furthermore, global harmonic potentials have served as a useful starting point for modeling many molecules and have been used with success to simulate vibronic spectra.[6, 56, 57, 19]
@@ III.1 Two-dimensional harmonic potentials
To demonstrate the Hagedorn approach in the simplest multi-dimensional cases, we start with two-dimensional models so that “exact” quantum calculations are readily accessible for the verification of the results.
@@
We assume that the initial, excited-state electronic surface $V_{e}$ and the final, ground-state electronic surface $V_{g}$ can both be described by a quadratic function (11). The ground-state surface $V_{g}$ is centered around zero, i.e., $q_{\text{ref},g}=(0,0)$. In all cases, the initial wavepacket is centered at $q_{0}=q_{\text{ref},e}=(15,-15)$, i.e. at the minimum of the excited-state surface $V_{e}$, and has zero momentum $p_{0}=(0,0)$.
@@
Keeping $V_{e}$ the same in all cases, we then construct three different ground-state surfaces $V_{g}$ to demonstrate the effects of mode displacement, distortion, and Duschinsky rotation in a cumulative fashion. In each case, a Gaussian wavepacket corresponding to the ground vibrational state of $V_{e}$ is propagated for 40000 a.u. (20000 steps with a time step of 2 a.u.) on the surface $V_{g}$. For each system, we provide examples of emissions from various vibrational levels, including the ground vibrational level, a singly excited level, a level with higher excitations in one mode, and two levels with multi-mode excitations. Specifically, we simulate SVL spectra from initial vibrational levels $1^{a}2^{b}$, $(a,b)\in\{(0,0),(1,0),(0,3),(1,1)(2,1)\}$, where $a,b$ denote the vibrational quantum numbers in modes 1 and 2 of the excited electronic state.
@@
The autocorrelation function of the Hagedorn wavepacket associated with each initial vibrational level is computed every two steps using the overlap expressions derived in Ref. 49. The SVL emission spectra are then evaluated by Fourier transforming the autocorrelation functions multiplied with a Gaussian damping function, which results in a spectral broadening with a half-width at half-maximum of $50\,\text{cm}^{-1}$.
@@
Figure 1 contains the SVL emission spectra of the three systems computed using Hagedorn wavepackets. The spectral intensities in all spectra are scaled by the highest peak of the $1^{0}2^{0}$ spectrum in the displaced, distorted and rotated system, which can be considered as the most comprehensive harmonic description of a real, anharmonic system. The other two systems can be then thought of as simplified approximations of the full harmonic system. Each spectrum is shifted so that the transition to the vibrational ground state ($1^{a}_{0}2^{b}_{0}$, where the subscripts denote the final vibrational quantum numbers in the ground electronic state after the transition) is at $0\,\text{cm}^{-1}$.
@@
If the ground and excited electronic surfaces no longer have the same corresponding vibrational frequencies, the so-called “mode distortion” leads to changes in peak positions as well as in intensity (second column in Fig. 1) in addition to the effects from displacement. In the displaced and distorted system, the initial wavepackets are kept the same, but $V_{g}$ now has a diagonal Hessian matrix $\kappa_{g,\text{distorted}}$ corresponding to vibrational wavenumbers $\tilde{\nu}_{1}^{\prime\prime}=750\,\text{cm}^{-1}$ and $\tilde{\nu}_{2}^{\prime\prime}=2200\,\text{cm}^{-1}$. As the ground electronic surface changes, the possible peak positions adapt accordingly at the combination of new frequencies, which can be observed by comparing the ordering and positions of the labeled peaks on the $1^{0}2^{0}$ spectra in the first and second columns. In this particular case, since $\tilde{\nu}_{2}^{\prime\prime}\approx 3\tilde{\nu}_{1}^{\prime\prime}$, the spectral peaks exhibit a more regular spacing, and certain broadened peaks, e.g. for $1^{0}_{3}2^{0}_{0}$ and $1^{0}_{0}2^{0}_{1}$ (labeled together on the $1^{0}2^{0}$ spectrum), actually overlap in this displaced and distorted system. The mode distortion also alters the intensity pattern, which is particularly evident in the $1^{1}2^{1}$ and $1^{2}2^{1}$ spectra (compare the first and second columns).
Finally, in the Duschinsky-rotated system (third column in Fig. 1), $V_{g}$ is modified so that $\kappa_{g,\text{rotated}}=R({{20}^{\circ}})\cdot\kappa_{g,\text{distorted}}$, where
@@
is a rotation matrix. The analysis of spectra here becomes more challenging, even in two-dimensional systems. We observe significant differences in the intensity pattern compared to unrotated systems. In our particular case, the frequency range of the spectral region also appears to be decreased. From a time-independent perspective, this could be explained by the fact that the Duschinsky rotation reduces the overlaps or Franck-Condon factors between the initial wavepacket and higher vibrational states on the ground electronic surface. It is worth noting that in this most general harmonic model, the Hagedorn wavepackets do not maintain the separable form of (10) during propagation, since $Q_{t}$ is no longer diagonal when evolved under a Duschinsky-rotated potential. With Hagedorn’s ingenious construction, the propagation algorithm maintains its simplicity even for general, non-separable Hagedorn wavepackets, which are not simply tensor products of eigenfunctions of one-dimensional harmonic oscillators.
In order to validate the Hagedorn approach, we also propagated the initial states using the second-order Fourier split-operator algorithm with $256\times 256$ grid points; however, unlike the Hagedorn approach, the quantum dynamics must be performed separately for each initial vibrational level in each system. Where applicable, the spectra were also evaluated using the TGA[23, 50, 25] (for $1^{0}2^{0}$) and its extended version[31, 14] (for $1^{1}2^{0}$ and $1^{1}2^{1}$). We compare the spectra obtained using Hagedorn wavepackets to the results from “exact” quantum calculations as well as (extended) TGA simulations in Fig. 2. Before calculating the differences, the spectra were all scaled and shifted as in Fig. 1. As expected, we observe excellent agreement between the methods. The comparison with the (extended) TGA calculations, which are performed in Heller’s parametrization[23, 51] of the Gaussian, which is substantially different from Hagedorn’s, serves as a further validation. In addition, ETGA employs a different approach to propagate the linear polynomial factor.
@@
A notable advantage of Hagedorn wavepackets over the ETGA and the previously reported generating function approach[19] is their ability to simulate dynamics starting from arbitrary initial vibrational levels, going beyond linear polynomial factors. Among the examples shown, the ETGA approach can compute emission from the $1^{1}2^{1}$ vibrational level only for cases without Duschinsky rotation, where the system can be separated into two one-dimensional systems each with at most a single excitation. The $1^{1}2^{1}$ spectra are then computed using the correlation function obtained by multiplying the correlation functions from the two one-dimensional systems.

### v2

@@ Abstract
In single vibronic level (SVL) fluorescence experiments, the electronically excited initial state is also excited in one or several vibrational modes. Whereas computing all contributing Franck-Condon factors individually becomes impractical in large systems, a time-dependent formalism has not been applied to simulate emission from arbitrary initial vibrational levels. Here, we use Hagedorn functions, which are products of a Gaussian and carefully generated polynomials, to represent SVL initial states. In systems where the potential is at most quadratic, the Hagedorn functions are exact solutions to the time-dependent Schrödinger equation and can be propagated with the same equations of motion as a simple Gaussian wavepacket. Having developed an efficient recursive algorithm to compute the overlaps between two Hagedorn wavepackets, we can now evaluate emission spectra from arbitrary vibronic levels using a single trajectory. We validate the method in two-dimensional global harmonic models by comparing it with quantum split-operator calculations. Additionally, we study the effects of displacement, distortion (squeezing), and Duschinsky rotation on SVL spectra. Finally, we demonstrate the applicability of the Hagedorn approach to high-dimensional systems on an example of displaced, distorted, and Duschinsky-rotated harmonic model with 100 degrees of freedom.
@@ I Introduction
Vibronic spectroscopy provides valuable insights into the quantum dynamics of molecules.[1] In single vibronic level (SVL) fluorescence spectrocopy, tuned narrow-band lasers are used to prepare a population of molecules excited to a specific vibronic level. The subsequent fluorescence decay of such a system can then be measured and provide important information about vibrational structures, excited-state relaxation, as well as the vibronic coupling between ground and excited states.[2, 3, 4, 5, 6, 7, 8]
@@
The simulation of SVL spectra of small molecules has been carried out using time-independent sum-over-state expressions,[9, 10, 11, 12, 13, 14, 15] but the computation of Franck-Condon factors in larger molecules becomes more challenging with the increasing number of vibrational states involved, especially due to the non-separable overlap integrals under Duschinsky mixing effects. A time-dependent approach is more efficient for larger systems and constitutes a more fundamental method that is not only naturally suited to simulate time-resolved and multidimensional spectroscopy[16, 17, 18, 19, 20] but can also readily incorporate Duschinsky rotation, Herzberg-Teller,[21, 22, 23, 24, 25, 26] and temperature effects.[27, 28, 29, 30] While time-dependent approaches have been applied to SVL[31, 32] and similar experiments,[33, 34, 35] their practical application has so far been confined to initial states that are excited at most to the first excited vibrational level and only in a single vibrational mode.
@@
For absorption or emission from the ground vibrational state, the thawed Gaussian approximation[36, 37, 38] (TGA) has been successfully applied to simulate vibrationally resolved electronic spectra in situations where a single Gaussian wavefunction is sufficient to adequately describe the dynamics.[39, 40, 41, 28, 18, 29, 42, 43] However, the SVL emission process involving a vibrationally excited initial state cannot be described by a single Gaussian wavepacket even at the initial time. Within the harmonic approximation, it may instead be described, at all times, by a Gaussian multiplied by a polynomial. Tapavicza’s efficient generating function approach[32] can treat initial states singly excited in one mode, whereas the extended TGA[44, 22, 28, 25, 45, 46] (ETGA) may accommodate simultaneous single excitations in multiple modes, but only in systems without Duschinsky rotation.[47] A more general and robust method is needed to address situations where the simultaneous excitation of multiple vibrational modes to higher levels is involved.
@@
By generalizing Dirac’s ladder operators for one-dimensional quantum harmonic oscillators to arbitrary dimensions, Hagedorn introduced a special pair of raising and lowering operators that generate a series of functions in the form of a Gaussian wavepacket multiplied by a polynomial.[48, 49, 50, 51, 52, 53] Unlike simple direct products of one-dimensional Hermite functions, Hagedorn functions are exact solutions to the time-dependent Schrödinger equation (TDSE) with an arbitrary many-dimensional harmonic potential when the parameters of the associated Gaussian are propagated with classical-like equations of motion. The Hagedorn functions also form a complete orthonormal basis that may be used to expand arbitrary numerically exact solutions to the TDSE with a general potential.[54] However, in this work, we consider exclusively Hagedorn wavepackets consisting of only one such Hagedorn function at all times. Henceforth, we will use “Hagedorn wavepackets” and “Hagedorn functions” synonymously.
@@
Despite their promising properties[52, 55, 56, 57, 53] and a few applications in quantum and chemical physics,[58, 59, 54, 60, 61, 62, 63, 64, 65, 66] Hagedorn wavepackets have yet to be applied to spectroscopic simulation, since one important ingredient, an efficient way to compute the correlation functions between two Hagedorn wavepackets, was missing. Having developed an efficient algebraic algorithm for this overlap in Ref. 67, we can now apply the Hagedorn approach to the simulation of SVL spectra.
@@
Here, we use the Hagedorn wavepacket dynamics to compute SVL emissions from different vibrational levels in two-dimensional harmonic models in order to validate the results against “exact” quantum split-operator calculations and, where applicable, (extended) TGA simulations. We also examine how nuclear displacement, mode distortion (squeezing), and Duschinsky rotation influence SVL spectra by analyzing the different spectra arising from three harmonic models that gradually incorporate these effects. One important advantage of our approach is that the propagation of Hagedorn wavepackets does not incur any additional computational cost in harmonic potentials beyond that of propagating a single Gaussian wavepacket. The Hagedorn approach is thus suitable for much higher dimensions than grid-based quantum methods. As a demonstration, we compute spectra in a two-state harmonic model system with 100 vibrational degrees of freedom.
@@ II Theory
Let $|K\rangle\equiv|{e,K}\rangle$ denote a specific vibrational state on the adiabatic potential energy surface of the excited electronic state $e$, where $K=(K_{1},...,K_{D})$ is the multi-index of non-negative integers specifying the vibrational quantum numbers in the $D$ normal modes. Within the Condon approximation, the emission rate from $|K\rangle$ to the ground electronic state $g$ may be obtained by the Fourier transform[68, 32]
@@
Solving the time-dependent Schrödinger equation is a challenging numerical problem. Hagedorn’s semiclassical wavepackets offer a practical approach to evolve a wavepacket in higher-dimensional systems.[48, 51, 53] In Hagedorn’s parametrization,[54, 53, 69] a normalized $D$-dimensional Gaussian wavepacket is written as
$\varphi_{0}[\Lambda_{t},S_{t}](q)=\frac{1}{(\pi\hbar)^{D/4}\sqrt{\det(Q_{t})}}\\ \times\exp\left(\frac{i}{\hbar}\left(\frac{1}{2}x^{T}\cdot P_{t}\cdot Q_{t}^{-1}\cdot x+p_{t}^{T}\cdot x+S_{t}\right)\right],$ (3)
with the shifted position $x:=q-q_{t}$, a phase factor $S_{t}$ related to the classical action, and a set of time-dependent parameters $\Lambda_{t}=(q_{t},p_{t},Q_{t},P_{t})$, where $q_{t}$ and $p_{t}$ represent the position and momentum of the wavepacket’s center, and $Q_{t}$ and $P_{t}$ are complex-valued $D$-dimensional matrices related to the position and momentum covariances.[69] The two matrices replace the “width matrix” $A_{t}=P_{t}\cdot Q_{t}^{-1}$ used in Heller’s parametrization[70, 53, 69] and satisfy the symplecticity conditions [52, 53]
@@
Hagedorn defined a special pair of raising and lowering operators,[51, 53]
$A^{\dagger}$ $\equiv A^{\dagger}[\Lambda_{t}]:=\frac{i}{\sqrt{2\hbar}}\left(P_{t}^{\dagger}\cdot\hat{x}-Q_{t}^{\dagger}\cdot\hat{y}\right),$ (6)
$A$ $\equiv A[\Lambda_{t}]\hphantom{{}^{\dagger}}:=-\frac{i}{\sqrt{2\hbar}}\left(P_{t}^{T}\cdot\hat{x}-Q_{t}^{T}\cdot\hat{y}\right),$ (7)
@@
where $\langle j\rangle=(0,\dots,0,1,0,\dots,0)$ is the $j$-th unit vector in $D$ dimensions.[48, 51, 52, 53] In general, the polynomial factor cannot be expressed as a tensor product of univariate polynomials.[50, 71, 72, 53]
@@
which are the same as the equations for propagating a Gaussian wavepacket.[51, 53, 69] Here, the subscript $t$ denotes the time, $m$ the real symmetric mass matrix (scalar in mass-weighted normal-mode coordinates), and $L_{t}$ is the Lagrangian. Remarkably, with Hagedorn wavepackets, the SVL emission from any vibrational level can be evaluated from the same trajectory $(\Lambda_{t},S_{t})$ at zero cost beyond propagating a simple Gaussian.
To finally obtain SVL spectra from the dynamics results, the evaluation of overlap integrals (2) between two Hagedorn wavepackets with different Gaussian parameters is needed. However, a straightforward analytical expression, like that available for overlaps between two Gaussian wavepackets, has remained elusive for Hagedorn wavepackets. Past applications of Hagedorn wavepackets require numerical quadrature for integrals, which becomes challenging in higher dimensions.[54, 60, 62, 63, 64, 65, 66] In Ref. 67, we have proposed instead a recursive algebraic scheme to compute these overlaps. This allows us to apply Hagedorn wavepackets in much higher dimensions, paving the way for their spectroscopic applications. In real molecular applications, the cost of computing the overlaps becomes negligible compared to the costly polyatomic electronic structure calculations, which only need to be carried out once for the central Gaussian trajectory, and do not have to be repeated for the vibrationally excited initial states.
Hagedorn wavepackets can be seen as a generalization of coherent states,[73, 51, 55, 72, 74] and our approach is thus connected to the generating function formalism proposed by Huh and Berger.[31] Although their expressions, relying on multivariate Hermite polynomials, are also general for arbitrary initial vibrational levels, no implementation or application was given. As far as we know, the only practical formulation is that of Tapavicza,[32] which is so far limited to single excitation in one mode. The Hagedorn approach, in addition to its generality, also provides a more intuitive time-dependent picture and offers a straightforward propagation scheme.
In this study, we focus solely on global harmonic models so that the results can be analyzed without additional approximation or errors. Whereas in many situations anharmonicity[24, 75, 76, 77] and non-adiabatic[78] effects must be taken into account, adiabatic harmonic potentials have served as a useful starting point for computing and assigning the spectra of many molecules.[79, 80, 9, 81, 82, 83, 34, 32, 84, 85, 35, 86]
@@ III.1 Two-dimensional harmonic potentials
We begin by considering two-dimensional model systems. They are the simplest multi-dimensional cases where, in the presence of the Duschinsky rotation effects, the dynamics of Hagedorn wavepackets differ from that of simple products of one-dimensional Hermite functions. Additionally, numerically exact quantum calculations are readily accessible for the verification of the surprisingly complicated spectra of these models.
@@
We assume that the initial, excited-state electronic surface $V_{e}$ and the final, ground-state electronic surface $V_{g}$ can both be described by a quadratic function (11). In all cases, the ground-state surface $V_{g}$ is centered at $q_{\text{ref},g}=(0,0)$., and the initial wavepacket is centered at $q_{0}=q_{\text{ref},e}=(15,-15)$, i.e., at the minimum of the excited-state surface $V_{e}$, and has zero momentum $p_{0}=(0,0)$.
@@
Keeping $V_{e}$ the same in all systems, we then construct three different ground-state surfaces $V_{g}$ to demonstrate the effects of mode displacement, distortion, and Duschinsky rotation in a cumulative fashion. In each case, a Gaussian wavepacket, which gives rise to the emission spectrum from the ground vibrational level of $V_{e}$, is propagated for 40000 a.u. (20000 steps with a time step of 2 a.u.) on the surface $V_{g}$.
For each system, we provide examples of emissions from various vibrational levels, including the ground vibrational level, a singly excited level, a level with higher excitations in one mode, and two levels with multi-mode excitations. Specifically, we simulate SVL spectra from initial vibrational levels $1^{a}2^{b}$, $(a,b)\in\{(0,0),(1,0),(0,3),(1,1),(2,1)\}$, where $a,b$ denote the vibrational quantum numbers in modes 1 and 2 of the excited electronic state.
@@
The autocorrelation function of the Hagedorn wavepacket associated with each initial vibrational level is computed every two steps using the overlap expressions derived in Ref. 67. The SVL emission spectra are then evaluated by Fourier transforming the autocorrelation functions multiplied with a Gaussian damping function, which results in a spectral broadening with a half-width at half-maximum of $50\,\text{cm}^{-1}$. Each spectrum is shifted so that the transition to the vibrational ground state ($1^{a}_{0}2^{b}_{0}$, where the subscripts denote the final vibrational quantum numbers in the ground electronic state after the transition) is at $0\,\text{cm}^{-1}$.
Figure 1: Comparison of the $1^{2}2^{1}$ SVL spectra (upper panel) computed using quantum split-operator calculations (black line) and Hagedorn wavepacket dynamics (red dashes) in a displaced, distorted, and Duschinsky-rotated two-dimensional harmonic system, with the absolute difference between the two simulated spectra shown in blue in the lower panel.
In order to validate the Hagedorn approach, we also propagated the initial states using the second-order Fourier split-operator algorithm with $256\times 256$ grid points. Figure 1 shows the comparison between the exact quantum and the Hagedorn dynamics results in the most complicated case treated here, i.e., the SVL spectrum involving a higher mixed initial excitation ($1^{2}2^{1}$) in the displaced, distorted, and Duschinsky-rotated system, which can be considered as the most comprehensive harmonic description of a real, anharmonic system. The spectrum simulated from Hagedorn dynamics is not only visually indistinguishable from the grid-based quantum calculation but the absolute differences are also extremely small (on the scale of $<10^{-11}$). The excellent agreement between the Hagedorn and exact quantum results is also observed for the other SVL spectra simulated in all three systems, which we now discuss in detail.
@@
Figure 2 contains the SVL emission spectra of the three systems computed using Hagedorn wavepackets. The spectral intensities in all spectra are scaled by the highest peak of the $1^{0}2^{0}$ spectrum in the displaced, distorted and rotated system because the other two systems can be thought of as simplified approximations of this most general harmonic system.
@@
If the ground and excited electronic surfaces no longer have the same vibrational frequencies, the so-called “mode distortion” leads to changes in peak positions as well as in intensity (second column in Fig. 2) in addition to the effects from displacement. In the displaced and distorted system, the initial wavepackets are kept the same, but $V_{g}$ now has a different (although still diagonal) Hessian matrix $\kappa_{g,\text{distorted}}$ corresponding to vibrational wavenumbers $\tilde{\nu}_{1}^{\prime\prime}=750\,\text{cm}^{-1}$ and $\tilde{\nu}_{2}^{\prime\prime}=2200\,\text{cm}^{-1}$. As the ground electronic surface changes, the possible peak positions adapt accordingly to the combination of new frequencies, which can be observed by comparing the ordering and positions of the labeled peaks on the $1^{0}2^{0}$ spectra in the first and second columns. In this particular case, since $\tilde{\nu}_{2}^{\prime\prime}\approx 3\tilde{\nu}_{1}^{\prime\prime}$, the spectral peaks exhibit a more regular spacing, and certain broadened peaks, e.g. for $1^{0}_{3}2^{0}_{0}$ and $1^{0}_{0}2^{0}_{1}$ (labeled together on the $1^{0}2^{0}$ spectrum), actually overlap in this displaced and distorted system. The mode distortion also alters the intensity pattern, which is particularly evident in the $1^{1}2^{1}$ and $1^{2}2^{1}$ spectra (compare the first and second columns).
Finally, in the Duschinsky-rotated system (third column in Fig. 2), $V_{g}$ is modified so that $\kappa_{g,\text{rotated}}=R({{20}^{\circ}})^{T}\cdot\kappa_{g,\text{distorted}}\cdot R({{20}^{\circ}})$, where
@@
is a rotation matrix. The analysis of spectra here becomes more challenging, even in two-dimensional systems. We observe significant differences in the intensity pattern compared to unrotated systems. In our particular case, the frequency range of the spectral region also appears to be decreased. From a time-independent perspective, this could be explained by the fact that the Duschinsky rotation reduces the Franck-Condon overlaps between the initial wavepacket and higher vibrational states on the ground electronic surface. It is important to note that in this most general harmonic model, the Hagedorn wavepackets do not maintain the separable form of (10) during propagation, since $Q_{t}$ is no longer diagonal when evolved under a Duschinsky-rotated potential. With Hagedorn’s ingenious construction, the propagation algorithm maintains its simplicity even for general, non-separable Hagedorn wavepackets, which are not simply tensor products of eigenfunctions of one-dimensional harmonic oscillators.
All spectra shown in Fig. 2 were obtained with the Hagedorn approach and are validated against quantum split operator results in Fig. 3. In contrast to the Hagedorn approach, the quantum dynamics must be performed separately for each initial vibrational level in each system. Where applicable, the Hagedorn results are also compared to the spectra evaluated using the TGA[36, 68, 38] (for $1^{0}2^{0}$) and its extended version[44, 22] (for $1^{1}2^{0}$ and $1^{1}2^{1}$). Before calculating the differences, the spectra were all scaled and shifted as in Fig. 2. As expected, we observe excellent agreement between the methods. The comparison with the (extended) TGA results serves as a further validation, as these calculations are performed in Heller’s Gaussian parametrization[36, 69] that differs substantially from Hagedorn’s. In addition, ETGA employs a different approach to propagate the linear polynomial prefactor of the Gaussian.
@@
A notable advantage of Hagedorn wavepackets over the ETGA and the previously reported generating function approach[32] is their ability to simulate dynamics starting from arbitrary initial vibrational levels, going beyond linear polynomial factors. Among the examples shown, the ETGA approach can compute emission from the $1^{1}2^{1}$ vibrational level only for cases without Duschinsky rotation, where the system can be separated into two one-dimensional systems each with at most a single excitation. The $1^{1}2^{1}$ spectra were then computed by Fourier transforming the correlation function obtained by multiplying the correlation functions from the two one-dimensional systems.